### v2

# Reduced dynamics for models of pattern formation

Abstract.
Chemical and biochemical reactions can exhibit a wide range of complex behaviors, including multiple steady states, oscillatory patterns, and chaotic dynamics. These phenomena have captivated researchers for many decades. Notable examples of oscillating chemical systems include the Briggs–Rauscher, Belousov–Zhabotinskii, and Bray–Liebhafsky reactions, where periodic variations in concentration are often visualized through observable color changes. These systems are typically modeled by a set of partial differential equations coupled through nonlinear interactions.
Upon closer analysis, it appears that the dynamics of these chemical/biochemical reactions may be governed by only a finite number of spatial Fourier modes. We can also draw the same conclusion in fluid dynamics, where it has been shown that, over long periods, the fluid velocity is determined by a finite set of Fourier modes, referred to as determining modes. In this article, we introduce the concept of determining modes for a two-species chemical models, which covers models such as the Brusselator, the Gray-Scott model, and the Glycolysis model [1, 18]. We demonstrate that it is indeed sufficient to characterize the dynamic of the model using only a finite number of spatial Fourier modes.
Key words and phrases:
Reduced Dynamics, Galerkin Approximation, Asymptotic Analysis, Reaction-Diffusion Equation, Pattern Formation, Determining Modes
1991 Mathematics Subject Classification:
35Q92, 35B40, 35K57, 65N30
This work was supported by the Austrian Science Fund (FWF) 10.55776/P44681.
1. Introduction
Pattern formation arises from the interactions among different components, potentially influenced by their environment. Alan Turing, a cryptographer and pioneer in computer science, developed algorithms to describe complex patterns emerging from simple inputs and random fluctuations. In his seminal 1952 paper [19], Turing proposed that the interaction between two biochemical substances with differing diffusion rates could generate biological patterns. His work addressed a key question in morphogenesis: how a single egg can develop into a complex organism. In the mathematical framework, an activator protein enhances both its own production and the production of an inhibitory protein, which solely functions to inhibit the activator. It has been observed that a stable homogeneous pattern can become unstable if the inhibitor diffuses more rapidly than the activator. The interaction between the concentrations of these proteins leads to pattern formation, with their spatiotemporal evolution governed by coupled reaction-diffusion systems, known as the activator-inhibitor model. This phenomenon is referred to as diffusion-driven instability, or Turing instability. The fundamental phenomenon in activator-inhibitor systems is that a small deviation from spatial homogeneity can trigger strong positive feedback, amplifying the deviation further. Nonlinearities in the local dynamics, such as those introduced by the inhibitor concentration, can then saturate the Turing instability, resulting in a stable and spatially inhomogeneous pattern. For a more detailed discussion, refer to the recent works by Upadhyay and Iyengar [13], Wei and Winter [20], Keener [12], and Perthame [15].
When analyzing these patterns in spatial Fourier modes, it is observed that typically only a finite number of modes play significant roles, while higher modes can often be neglected due to their dissipation. This raises the questions: how many modes are necessary to effectively analyze and characterize the system’s dynamical behavior? Is it essential to approximate the solution with high fidelity, or is it sufficient to model the system using only a few modes?
The concept of determining modes becomes particularly relevant in this context. Originally introduced in the study of fluid dynamics and turbulence, the notion of determining modes is crucial for understanding complex phenomena like turbulence, which is characterized by chaotic and unpredictable fluid motion. Identifying and analyzing the underlying structures, or modes, within turbulent flows is essential for accurately predicting the behavior of fluid systems. Determining modes were introduced to identify the parameters that control turbulent flow; see, for example, the monograph by Foiaş [5, Chapter III] and the articles by Titi [10] and Foiaş and Temam [6]. This concept is practically significant. Modes are said “determining” if the high modes are asymptotically controlled by these lower modes. By identifying these determining modes, one can characterize the dominant patterns or structures within the flow. In the long run, a finite number of parameters — the coefficients of the determining modes — can effectively describe the entire system.
The concept of determining modes has been successfully applied in numerical simulations to reduce computational time. It is also used in data assimilation [21] and control [11]. In the context of reaction-diffusion equations, determining modes have been discussed in [3], where the systems considered are of a monotone type.
Let $D\coloneqq[0,1]^{2}$. In this paper, we are interested to the behavior in the limit $t\to\infty$ of a two-species model of the form
(1.1) $\partial_{t}{u}(t)&=d_{1}\Delta u(t)+a_{1}u(t)+b_{1}v(t)-\gamma u(t)v^{2}(t)+g_{1}(t),&\mbox{ in }D,\\ \partial_{t}{v}(t)&=d_{2}\Delta v(t)+a_{2}u(t)+b_{2}v(t)+\gamma u(t)v^{2}(t)+g_{2}(t),&\mbox{ in }D,\\ u(0)&=u_{0}\geq 0,\;v(0)=v_{0}\geq 0,&\mbox{ in }D,\end{aligned}\right.$
where for $j=1,2$, $d_{j},a_{j},b_{j},\gamma$ are constants and $g_{j}$ are time-dependant maps to be specified later. We will consider the problem associated with zero Neumann boundary conditions on $\partial D$.
Concerning the solution theory, we can refer to [16], where the author discusses questions related to the global solution to a class of reaction-diffusion systems that satisfies two major properties: the quasi-positivity (P) and the mass-control (M). With these two properties valid, it is sometime possible to show the global existence of the classical solution. Moreover, the solution remains positive for all time. That applies, for instance, to the Brusselator model [7], which is a particular case of 1.1, where $a_{1}=a_{2}=0$; $b_{1}>0,b_{2}=b_{1}+1$; $\gamma=1$; $g_{1}=0,g_{2}>0$. The Brusselator satisfies the properties (P) + (M), and it was shown, see [8, Proposition 2], [17, Theorem 1, Page 140], that it has a unique global classical solution.
The system 1.1 fits also within the framework established by McGough and Riley in [14], where they develop a general class of reaction-diffusion problems characterized by telescoping nonlinearities, which arise in contexts such as chemical kinetics and population dynamics. These problems have special properties, which for the case of the two-species model 1.1 are given as follows:
  1. ($P_{1})$
    $d_{1},d_{2}>0$,
  2. ($P_{2})$
    $g_{1},g_{2}\in L^{\infty}(0,\infty;L^{p}(D))$, for any $p\geq 2$,
  3. ($P_{3})$
    $\gamma\geq 0$,
  4. ($P_{4})$
    For all $u,v\geq 0$, $b_{1}v+g_{1}\geq 0\mbox{ and }a_{2}u+g_{2}\geq 0,$
  5. ($P_{5})$
    There exist constants $c_{1}\geq 0$ and $c_{2}>0$ such that $a_{1}u+b_{1}v-\gamma uv^{2}+g_{1}\leq c_{1}(1-c_{2}u)$,
  6. ($P_{6})$
    $a_{1}+a_{2}\leq 0$ and $b_{1}+b_{2}\leq 0$.
If the properties ($P_{1})$-($P_{6})$ hold, McGough and Riley can demonstrate several theoretical results, including that all positive solutions to the system remain bounded and smooth for all time. This result holds for a domain $D\subset\mathbb{R}^{N}$, $N=2,3$, with a piecewise smooth boundary.
The goal of this paper is to demonstrate that the asymptotic behavior of the system of partial differential equations given by 1.1 can be effectively monitored by projecting the system onto a finite-dimensional space, known as determining modes. For a precise definition of the number of determining modes, we refer to Definition 2.3. In fluid dynamics, the number of determining modes is often expressed in terms of the Grashof number. This number quantifies the ratio of buoyancy forces to viscous forces in a fluid, making it an effective parameter for assessing the balance of forces within the system. For more details, see [5, Section II.13]. In our work, we will adopt the same approach, characterizing the number of determining modes using the Grashof number. To be more precise, we estimate the number of determining modes using a nondimensional constant, $\boldsymbol{\mathrm{Gr}}$, known as the Grashof number.
The rest of the paper is organized as follows. In Section 2, we collect background materials and necessary assumptions on the data and present the main result. In Section 3, we introduce some necessary Lemmata and propositions for the proof of the main statement. In Section 4, we prove that 1.1 admits a set of determining modes, see Definition 2.3.
2. Background materials and the main statement
In this section, we will introduce standard notations and recall results on the solution theory of systems like 1.1. Then, we present the definition of determining modes and the main result. The proof of the main result is postponed to Section 4.
[…]
$\|u\|_{L^{p}}^{p}\coloneqq\int_{D}|u(x)|^{p}\,\mathrm{d}x,\,1\leq p<\infty,\quad\mbox{ and }\quad\|u\|_{L^{\infty}}\coloneqq\sup_{x\in D}|u(x)|.$
In the paper, $(u,v)$ is understood as the global smooth solution to 1.1, which existence is made precise in the following Lemmata, see also [14, Theorem 10].
Lemma 2.1.
We assume that the properties ($P_{1})$-($P_{6})$ hold for the system 1.1. If $u_{0},v_{0}\in L^{2}(D)$ then for the solution to the system 1.1 it holds that $u(t),v(t)\geq 0$ and $u(t),v(t)\in C^{2}(D)$ for all $t\geq 0$.
More generally, the following estimates hold for $(u,v)$, see e.g. [14, Theorem 2].
Lemma 2.2.
If $u_{0},v_{0}\in L^{p}(D)$, then $u,v\in L^{\infty}(0,\infty;L^{p}(D))$ for all $p\geq 2$.
2.1. Determining modes
Let $(\tilde{g}_{1},\tilde{g}_{2})$ be a deterministic perturbation of $(g_{1},g_{2})$ such that $(\tilde{g}_{1},\tilde{g}_{2})$ and $(g_{1},g_{2})$ have the same asymptotic behavior for large times, i.e.
(2.1) $\limsup_{t\to\infty}\|(g_{1}-\tilde{g}_{1})(t)\|_{L^{p}}+\limsup_{t\to\infty}\|(g_{2}-\tilde{g}_{2})(t)\|_{L^{p}}=0.$
To $(\tilde{g}_{1},\tilde{g}_{2})$ we associate the couple $(\tilde{u},\tilde{v})$ that satisfies for all $t\geq 0$ to the system
(2.2) $\partial_{t}{\tilde{u}}(t)&=d_{1}\Delta\tilde{u}(t)+a_{1}\tilde{u}(t)+b_{1}\tilde{v}(t)-\gamma\tilde{u}(t)\tilde{v}^{2}(t)+\tilde{g}_{1}(t),&\mbox{ in }D,\\ \partial_{t}{\tilde{v}}(t)&=d_{2}\Delta\tilde{v}(t)+a_{2}\tilde{u}(t)+b_{2}\tilde{v}(t)+\gamma\tilde{u}(t)\tilde{v}^{2}(t)+\tilde{g}_{2}(t),&\mbox{ in }D,\\ \tilde{u}(0)&=u_{0}\geq 0,\;\tilde{v}(0)=v_{0}\geq 0,&\mbox{ in }D,\end{aligned}\right.$
with zero Neumann boundary conditions on $\partial D$. We assume that the following properties holds for the system 2.2
  1. ($\tilde{P}_{1})$
    $d_{1},d_{2}>0$,
  2. ($\tilde{P}_{2})$
    $\tilde{g}_{1},\tilde{g}_{2}\in L^{\infty}(0,\infty;L^{p}(D))$, for any $p\geq 2$,
  3. ($\tilde{P}_{3})$
    $\gamma\geq 0$,
  4. ($\tilde{P}_{4})$
    For all $u,v\geq 0$, $b_{1}v+\tilde{g}_{1}\geq 0\mbox{ and }a_{2}u+\tilde{g}_{2}\geq 0,$
  5. ($\tilde{P}_{5})$
    There exist constants $\tilde{c}_{1}\geq 0$ and $\tilde{c}_{2}>0$ such that $a_{1}u+b_{1}v-\gamma uv^{2}+\tilde{g}_{1}\leq\tilde{c}_{1}(1-\tilde{c}_{2}u)$,
  6. ($\tilde{P}_{6})$
    $a_{1}+a_{2}\leq 0$ and $b_{1}+b_{2}\leq 0$.
In what follows, $(\tilde{u},\tilde{v})$ is understood as the global smooth solution to 2.2. Lemma 2.1 and Lemma2.2 are also valid for the system 2.2 provided that the properties ($\tilde{P}_{1})$-($\tilde{P}_{6})$ hold.
Let $\{\psi_{m}:m\in\mathbb{N}\}$ be the complete set of eigenfunctions of the Laplacian and $\{\lambda_{m}:m\in\mathbb{N}\}$ be the associated eigenvalues such that $-\Delta\psi_{m}=\lambda_{m}\psi_{m},\;m\in\mathbb{N}$. Since $\Delta$ is selfadjoint, $\{\psi_{m}:m\in\mathbb{N}\}$ is an orthogonal basis of $L^{2}(D)$. Moreover, we suppose that
$0<\lambda_{1}\leq\lambda_{2}\leq\cdots\leq\lambda_{m}\leq\cdots,\mbox{ and }\lambda_{m}\to\infty\mbox{ when }m\to\infty.$
Given some $M\in\mathbb{N}$, we define the finite subspace $\mathrm{H}_{M}\coloneqq\mbox{span}\big{\{}\psi_{m}:m\in\mathbb{N},\,m\leq M\big{\}}$ and denote the orthogonal projection of $L^{2}(D)$ onto $\mathrm{H}_{M}$ by $P_{M}$, the natural embedding by $I$, and define $Q_{M}$ by $Q_{M}\coloneqq I-P_{M}$. Given two indices $M$ and $N$ and the solution $(u,v)$ to 1.1, we define its Galerkin approximation by $(P_{M}u,P_{N}v)$.
We introduce the following definition of determining modes:
Definition 2.3.
The first modes $H_{M}:=\big{\{}\psi_{m}:m\in\mathbb{N},\,m\leq M\big{\}}$ and $H_{N}:=\big{\{}\psi_{m}:m\in\mathbb{N},\,m\leq N\big{\}}$ are called determining modes for the solution $(u,v)$ to 1.1, if
$\lim_{t\to\infty}\|P_{M}(u-\tilde{u})(t)\|_{L^{4}}+\lim_{t\to\infty}\|P_{N}(v-\tilde{v})(t)\|_{L^{4}}=0,$
implies that
$\lim_{t\to\infty}\|Q_{M}(u-\tilde{u})(t)\|_{L^{2}}+\lim_{t\to\infty}\|Q_{N}(v-\tilde{v})(t)\|_{L^{2}}=0,$
where, $(\tilde{u},\tilde{v})$, denote the solution to 2.2.
2.2. The main statement
We set $\xi\coloneqq u-\tilde{u}$, $\eta\coloneqq v-\tilde{v}$, $h_{1}\coloneqq g_{1}-\tilde{g}_{1}$, and $h_{2}\coloneqq g_{2}-\tilde{g}_{2}$. We denote
$g_{1}^{*}\coloneqq\limsup_{t\to\infty}\|g_{1}(t)\|_{L^{2}}^{2},\quad g_{2}^{*}\coloneqq\limsup_{t\to\infty}\|g_{2}(t)\|_{L^{2}}^{2}.$
We set $F\coloneqq c_{1}+g_{1}^{*}+g_{2}^{*}+A_{2}^{2}+B_{2}$ and $d\coloneqq\min\{d_{1},d_{2}\}$, with $B_{2}\coloneqq c_{1}|D|^{1/2}$. We denote the generalized Grashof number that corresponds to the system 1.1 by
$\boldsymbol{\mathrm{Gr}}\coloneqq\frac{F}{d^{2}\lambda_{1}}.$
Theorem 2.4.
We assume that the properties ($P_{1})$-($P_{6})$ hold for the system 1.1. If there exist $M,N\in\mathbb{N}$ such that
$\frac{\lambda_{M+1}+\lambda_{N+1}}{\lambda_{1}}>2\gamma^{2}d^{2}\boldsymbol{\mathrm{Gr}}^{4}+\boldsymbol{\mathrm{Gr}},$
then the first $(M,N)$ modes are determining for the solution $(u,v)$ to the system 1.1.
We postpone the proof of Theorem 2.4 to Section 4. In the first place, we show some preliminary estimates, notably Lemmas 3.1 , 3.2 and 3.3 that will be used as intermediate steps in proving Theorem 2.4.
3. Preliminaries
We will start with a Grownwall-type Lemma.
Lemma 3.1.
Let $\alpha$ be a locally integrable real valued function on $(0,\infty)$, satisfying for some $0<T<\infty$ the following conditions:
(3.1) $\gamma\coloneqq\liminf_{t\to\infty}\frac{1}{T}\int_{t}^{t+T}\alpha(s)\,\mathrm{d}s>0,\quad\Gamma\coloneqq\limsup_{t\to\infty}\frac{1}{T}\int_{t}^{t+T}\alpha^{-}(s)\,\mathrm{d}s<\infty,$
where $\alpha^{-}(t)=\max\{-\alpha(t),0\}$. Further let $\beta$ be a real-valued measurable function defined on $(0,\infty)$ such that
(3.2) $\lim_{t\to\infty}\beta(t)=0.$
Suppose that $X$ is an absolutely continuous non-negative function on $(0,\infty)$, with $X_{0}\coloneqq X(0)$, such that
[…]
Then $X(t)\to 0$ as $t\to\infty$. Actually, we have
$X(t)\leq X_{0}\Gamma^{\prime}\mathrm{e}^{-(\gamma/2T)(t-0)}+\frac{2\Gamma^{\prime}T}{\gamma}\sup_{0\leq\tau\leq t}\|\beta(\tau)\|_{L^{2}},$
with $\Gamma^{\prime}\coloneqq\mathrm{e}^{\Gamma+1+\gamma/2}$.
Proof.
We deduce the results from the Gronwall lemma. See [4, Lemma 4.1] and [9, Lemma 4.1]. ∎
Lemma 3.2.
We assume that the properties ($P_{1})$-($P_{6})$ (resp. ($\tilde{P}_{1})$-($\tilde{P}_{6})$) hold for the system 1.1 (resp. 2.2. We set
$K_{M,N}(t)\coloneqq{}$ ${-\big{\langle}Q_{M}\xi(t)\big{(}a_{1}-\gamma v^{2}(t)\big{)}+Q_{N}\eta(t)\big{(}b_{1}-\gamma(v(t)+\tilde{v}(t))\tilde{u}(t)\big{)},Q_{M}\xi(t)\big{\rangle}}$
$-{\big{\langle}Q_{M}\xi(t)\big{(}a_{2}+\gamma v^{2}(t)\big{)}+Q_{N}\eta(t)\big{(}b_{2}+\gamma(v(t)+\tilde{v}(t))\tilde{u}(t)\big{)},Q_{N}\eta(t)\big{\rangle}}.$
If there exist $T>0$ and numbers $M,N\in\mathbb{N}$ such that
(3.3) $\begin{split}&\liminf_{t\to\infty}\bigg{(}\frac{1}{T}\int_{t}^{t+T}2d\bigg{(}\frac{\|\nabla Q_{M}\xi(s)\|_{L^{2}}+\|\nabla Q_{N}\eta(s)\|_{L^{2}}}{\|Q_{M}\xi(s)\|_{L^{2}}+\|Q_{N}\eta(s)\|_{L^{2}}}\bigg{)}^{2}\\ &\hskip 100.0pt+\frac{|K_{M,N}(t)|}{\big{(}\|Q_{M}\xi(s)\|_{L^{2}}+\|Q_{N}\eta(s)\|_{L^{2}}\big{)}^{2}}\,\mathrm{d}s\bigg{)}>0\end{split}$
and
(3.4) $\lim_{t\to\infty}\;\|P_{M}\xi(t)\|_{L^{4}}+\lim_{t\to\infty}\;\|P_{N}\eta(t)\|_{L^{4}}=0,$
then it holds that
$\lim_{t\to\infty}\;\|Q_{M}\xi(t)\|_{L^{2}}+\lim_{t\to\infty}\;\|Q_{N}\eta(t)\|_{L^{2}}=0.$
Proof.
We fix $M,N\in\mathbb{N}$, $t>0$, and $x\in D$. Standard calculations give
(3.5) $\partial_{t}\xi(t,x)-d_{1}\Delta\xi(t,x)$ $=\gamma\big{(}\tilde{u}(t,x)\tilde{v}^{2}(t,x)-u(t,x)v^{2}(t,x)\big{)}+a_{1}\xi(t,x)+b_{1}\eta(t,x)+h_{1}(t,x),$
(3.6) $\partial_{t}\eta(t,x)-d_{2}\Delta\eta(t,x)$ $=\gamma\big{(}u(t,x)v^{2}(t,x)-\tilde{u}(t,x)\tilde{v}^{2}(t,x)\big{)}+a_{2}\xi(t,x)+b_{2}\eta(t,x)+h_{2}(t,x).$
Observe the following calculations
$\gamma(uv^{2}$ $-\tilde{u}\tilde{v}^{2})+a_{2}\xi+b_{2}\eta$
$=(u-\tilde{u})\gamma v^{2}+(v+\tilde{v})\tilde{u}\gamma(v-\tilde{v})+a_{2}\xi+b_{2}\eta$
$=\xi\gamma v^{2}+(v+\tilde{v})\tilde{u}\gamma\eta+a_{2}\xi+b_{2}\eta$
$=(Q_{M}\xi)(a_{2}+\gamma v^{2})+\big{(}b_{2}+\gamma(v+\tilde{v})\tilde{u}\big{)}(Q_{N}\eta)+(P_{M}\xi)(a_{2}+\gamma v^{2})+\big{(}b_{2}+\gamma(v+\tilde{v})\tilde{u}\big{)}(P_{N}\eta).$
We take the scalar product in $L^{2}(D)$ with $Q_{M}\xi$ and $Q_{N}\eta$ on both sides of 3.53.6, respectively. It follows that
$\frac{1}{2}\frac{\,\mathrm{d}}{\,\mathrm{d}t}\Big{[}\|Q_{M}\xi(t)\|_{L^{2}}^{2}+\|Q_{N}\eta(t)\|_{L^{2}}^{2}\Big{]}+d_{1}\|\nabla Q_{M}\xi(t)\|_{L^{2}}^{2}+d_{2}\|\nabla Q_{N}\eta(t)\|_{L^{2}}^{2}+K_{M,N}(t)$
$=R_{M,N}(t)+\big{\langle}h_{1}(t),Q_{M}\xi(t)\big{\rangle}+\big{\langle}h_{2}(t),Q_{N}\eta(t)\big{\rangle}$
with
$K_{M,N}(t)\coloneqq{}$ ${-\big{\langle}Q_{M}\xi(t)\big{(}a_{1}-\gamma v^{2}(t)\big{)}+Q_{N}\eta(t)\big{(}b_{1}-\gamma(v(t)+\tilde{v}(t))\tilde{u}(t)\big{)},Q_{M}\xi(t)\big{\rangle}}$
$-{\big{\langle}Q_{M}\xi(t)\big{(}a_{2}+\gamma v^{2}(t)\big{)}+Q_{N}\eta(t)\big{(}b_{2}+\gamma(v(t)+\tilde{v}(t))\tilde{u}(t)\big{)},Q_{N}\eta(t)\big{\rangle}},$
$R_{M,N}(t)\coloneqq{}$ $\big{\langle}(P_{M}\xi(t))\big{(}a_{1}-\gamma v^{2}(t)\big{)}+\big{(}b_{1}-\gamma(v(t)+\tilde{v}(t))\tilde{u}(t)\big{)}(P_{N}\eta(t)),Q_{M}\xi(t)\big{\rangle}$
$+\big{\langle}(P_{M}\xi(t))\big{(}a_{2}+\gamma v^{2}(t)\big{)}+\big{(}b_{2}+\gamma(v(t)+\tilde{v}(t))\tilde{u}(t)\big{)}(P_{N}\eta(t)),Q_{N}\eta(t)\big{\rangle}.$
Next, we apply the Cauchy–Swcharz inequality, and for any $\varepsilon>0$, the Young inequality to get
$R_{M,N}(t)\leq\Big{(}\|P_{M}\xi(t)\|_{L^{4}}\|a_{1}-\gamma v^{2}(t)\|_{L^{4}}+\|P_{N}\eta(t)\|_{L^{4}}\|b_{1}-\gamma\big{(}v(t)+\tilde{v}(t)\big{)}\tilde{u}(t)\|_{L^{4}}\Big{)}\|Q_{M}\xi(t)\|_{L^{2}}$
$\qquad+\Big{(}\|P_{M}\xi(t)\|_{L^{4}}\|a_{2}+\gamma v^{2}(t)\|_{L^{4}}+\|P_{N}\eta(t)\|_{L^{4}}\|b_{2}+\gamma\big{(}v(t)+\tilde{v}(t)\big{)}\tilde{u}(t)\|_{L^{4}}\Big{)}\|Q_{N}\eta(t)\|_{L^{2}}$
$\leq\frac{1}{\varepsilon}\Big{(}\|P_{M}\xi(t)\|_{L^{4}}+\|P_{N}\eta(t)\|_{L^{4}}\Big{)}^{2}\Big{(}\gamma\|v(t)\|_{L^{8}}^{2}+\gamma\|\tilde{u}(t)\|_{L^{8}}^{2}+A_{4}\Big{)}^{2}+\varepsilon\Big{(}\|Q_{M}\xi(t)\|_{L^{2}}^{2}+\|Q_{N}\eta(t)\|_{L^{2}}^{2}\Big{)}$
$\leq\frac{6}{\varepsilon}\Big{(}\|P_{M}\xi(t)\|_{L^{4}}^{2}+\|P_{N}\eta(t)\|_{L^{4}}^{2}\Big{)}\Big{(}\gamma^{2}\|v(t)\|_{L^{8}}^{4}+\gamma^{2}\|\tilde{u}(t)\|^{4}_{L^{8}}+A_{4}^{2}\Big{)}+\varepsilon\big{(}\|Q_{M}\xi(t)\|_{L^{2}}+\|Q_{N}\eta(t)\|_{L^{2}}\big{)}^{2},$
with $A_{4}\coloneqq|D|^{1/4}\sum_{j=1}^{2}(|a_{j}|+|b_{j}|)$. Next, we set
$X(t)$ $\coloneqq\|Q_{M}\xi(t)\|_{L^{2}}^{2}+\|Q_{N}\eta(t)\|_{L^{2}}^{2},$
$\begin{split}\alpha(t)&\coloneqq 2d\bigg{(}\frac{\|\nabla Q_{M}\xi(t)\|_{L^{2}}+\|\nabla Q_{N}\eta(t)\|_{L^{2}}}{\|Q_{M}\xi(t)\|_{L^{2}}+\|Q_{N}\eta(t)\|_{L^{2}}}\bigg{)}^{2}+\frac{2|K_{M,N}(t)|}{\big{(}\|Q_{M}\xi(t)\|_{L^{2}}+\|Q_{N}\eta(t)\|_{L^{2}}\big{)}^{2}}-2\varepsilon,\end{split}$
$\beta(t)$ $\coloneqq\frac{6}{\varepsilon}\Big{(}\|P_{M}\xi(t)\|_{L^{4}}^{2}+\|P_{N}\eta(t)\|_{L^{4}}^{2}\Big{)}\Big{(}\gamma^{2}\|v(t)\|_{L^{4}}^{2}+\gamma^{2}\|\tilde{u}(t)\|^{2}_{L^{4}}+A_{4}^{2}\Big{)}+\frac{2}{\varepsilon}\|h_{1}(t)\|_{L^{2}}^{2}+\frac{2}{\varepsilon}\|h_{2}(t)\|_{L^{2}}^{2}.$
The results can be written in the following form
$\frac{\,\mathrm{d}}{\,\mathrm{d}t}X(t)+\alpha(t)X(t)\leq\beta(t).$
We show that $\beta$ (resp. $\alpha$) satisfies to the hypothesis 3.2 (resp. 3.1) of Lemma 3.1 and conclude.
The hypothesis 3.2 of Lemma 3.1 on $\beta$ holds thanks to the asymptotic behavior 2.1, Lemma 2.2, and 3.4.
The hypothesis 3.1 of Lemma 3.1 on $\alpha$ holds as a consequence of 3.3. Indeed, observe that if
$\liminf_{t\to\infty}\frac{1}{T}\int_{t}^{t+T}\alpha(s)\,\mathrm{d}s>2\varepsilon$
holds for some $T>0$ and $M,N\in\mathbb{N}$, then it implies that
(3.7) $\alpha(t)\geq 2d\bigg{(}\frac{\|\nabla Q_{M}\xi(t)\|_{L^{2}}+\|\nabla Q_{N}\eta(t)\|_{L^{2}}}{\|Q_{M}\xi(t)\|_{L^{2}}+\|Q_{N}\eta(t)\|_{L^{2}}}\bigg{)}^{2}+\frac{2|K_{M,N}(t)|}{\big{(}\|Q_{M}\xi(t)\|_{L^{2}}+\|Q_{N}\eta(t)\|_{L^{2}}\big{)}^{2}}-2\varepsilon.$
Before we continue with 3.7, we show that
(3.8) $\begin{split}&\frac{|K_{M,N}(t)|}{\big{(}\|Q_{M}\xi(t)\|_{L^{2}}+\|Q_{N}\eta(t)\|_{L^{2}}\big{)}^{2}}\\ &\leq{2\Big{(}\gamma^{2}\|v(t)\|_{L^{4}}^{4}+\gamma^{2}\|\tilde{u}(t)\|^{4}_{L^{4}}+A_{2}\Big{)}\bigg{(}\frac{\|\nabla Q_{M}\xi(t)\|_{L^{2}}+\|\nabla Q_{N}\eta(t)\|_{L^{2}}}{\|Q_{M}\xi(t)\|_{L^{2}}+\|Q_{N}\eta(t)\|_{L^{2}}}\bigg{)}},\end{split}$
with $A_{2}\coloneqq|D|^{1/2}\sum_{j=1}^{2}(|a_{j}|+|b_{j}|)$. By straightforward calculations, we get
$|K_{M,N}(t)|\leq{\|a_{1}-\gamma v^{2}(t)\|_{L^{2}}\|Q_{M}\xi(t)\|_{L^{4}}^{2}+\|b_{1}-\gamma\big{(}v(t)+\tilde{v}(t)\big{)}\tilde{u}(t)\|_{L^{2}}\|Q_{N}\eta(t)\|_{L^{4}}^{2}}$
$+{\Big{(}\|a_{2}+\gamma v^{2}(t)\|_{L^{2}}^{2}+\|b_{2}+\gamma\big{(}v(t)+\tilde{v}(t)\big{)}\tilde{u}(t)\|_{L^{2}}\Big{)}\|Q_{M}\xi(t)Q_{N}\eta(t)\|_{L^{2}}}$
$\leq\Big{(}\gamma\|v(t)\|_{L^{4}}^{2}+\gamma\|\big{(}v(t)+\tilde{v}(t)\big{)}\tilde{u}(t)\|_{L^{2}}+A_{2}\Big{)}\Big{(}{\|Q_{M}\xi(t)\|_{L^{4}}^{2}+\|Q_{N}\eta(t)\|_{L^{4}}^{2}}\Big{)}$
$+\Big{(}\gamma\|v(t)\|_{L^{4}}^{2}+\gamma\|\big{(}v(t)+\tilde{v}(t)\big{)}\tilde{u}(t)\|_{L^{2}}+A_{2}\Big{)}\Big{(}{\|Q_{M}\xi(t)Q_{N}\eta(t)\|_{L^{2}}}\Big{)}.$
Moreover, we have
${\|Q_{M}\xi(t)\|_{L^{4}}^{2}+\|Q_{N}\eta(t)\|_{L^{4}}^{2}}\leq{\|Q_{M}\xi(t)\|_{L^{2}}\|\nabla Q_{M}\xi(t)\|_{L^{2}}+\|Q_{N}\eta(t)\|_{L^{2}}\|\nabla Q_{N}\eta(t)\|_{L^{2}}}$
$\leq{\Big{(}\|Q_{M}\xi(t)\|_{L^{2}}+\|Q_{N}\eta(t)\|_{L^{2}}\Big{)}\Big{(}\|\nabla Q_{M}\xi(t)\|_{L^{2}}+\|\nabla Q_{N}\eta(t)\|_{L^{2}}\Big{)}},$
and
${\|Q_{M}\xi(t)Q_{N}\eta(t)\|_{L^{2}}}\leq{\|Q_{M}\xi(t)\|_{L^{4}}\|Q_{N}\eta(t)\|_{L^{4}}}$
$\leq{\frac{1}{2}\|Q_{M}\xi(t)\|_{L^{4}}^{2}+\frac{1}{2}\|Q_{N}\eta(t)\|_{L^{4}}^{2}}$
$\leq{\frac{1}{2}\|Q_{M}\xi(t)\|_{L^{2}}\|\nabla Q_{M}\xi(t)\|_{L^{2}}+\frac{1}{2}\|Q_{N}\eta(t)\|_{L^{2}}\|\nabla Q_{N}\eta(t)\|_{L^{2}}}$
$\leq{\frac{1}{2}\Big{(}\|Q_{M}\xi(t)\|_{L^{2}}+\|Q_{N}\eta(t)\|_{L^{2}}\Big{)}\Big{(}\|\nabla Q_{M}\xi(t)\|_{L^{2}}+\|\nabla Q_{N}\eta(t)\|_{L^{2}}\Big{)}}.$
Therefore, we arrive at
$\frac{|K_{M,N}(t)|}{\big{(}\|Q_{M}\xi(t)\|_{L^{2}}+\|Q_{N}\eta(t)\|_{L^{2}}\big{)}^{2}}$
$\leq{2\Big{(}\gamma\|v(t)\|_{L^{4}}^{2}+\gamma\|\big{(}v(t)+\tilde{v}(t)\big{)}\tilde{u}(t)\|_{L^{2}}+A_{2}\Big{)}\bigg{(}\frac{\|\nabla Q_{M}\xi(t)\|_{L^{2}}+\|\nabla Q_{N}\eta(t)\|_{L^{2}}}{\|Q_{M}\xi(t)\|_{L^{2}}+\|Q_{N}\eta(t)\|_{L^{2}}}\bigg{)}}$
$\leq{2\Big{(}\gamma\|v(t)\|_{L^{4}}^{2}+\gamma\|\tilde{u}(t)\|^{2}_{L^{4}}+A_{2}\Big{)}\bigg{(}\frac{\|\nabla Q_{M}\xi(t)\|_{L^{2}}+\|\nabla Q_{N}\eta(t)\|_{L^{2}}}{\|Q_{M}\xi(t)\|_{L^{2}}+\|Q_{N}\eta(t)\|_{L^{2}}}\bigg{)}}.$
The proof of 3.8 is complete.
We can continue with 3.7, where by using the estimate 3.8, we get
$\alpha(t)\geq 2d\bigg{(}\frac{\|\nabla Q_{M}\xi(t)\|_{L^{2}}+\|\nabla Q_{N}\eta(t)\|_{L^{2}}}{\|Q_{M}\xi(t)\|_{L^{2}}+\|Q_{N}\eta(t)\|_{L^{2}}}\bigg{)}^{2}$
$\hskip 50.0pt-{4\Big{(}\gamma\|v(t)\|_{L^{4}}^{2}+\gamma\|\tilde{u}(t)\|^{2}_{L^{4}}+A_{2}\Big{)}\bigg{(}\frac{\|\nabla Q_{M}\xi(t)\|_{L^{2}}+\|\nabla Q_{N}\eta(t)\|_{L^{2}}}{\|Q_{M}\xi(t)\|_{L^{2}}+\|Q_{N}\eta(t)\|_{L^{2}}}\bigg{)}}-2\varepsilon$
$\geq-{\frac{6}{d}\Big{(}\gamma^{2}\|v(t)\|_{L^{4}}^{4}+\gamma^{2}\|\tilde{u}(t)\|^{4}_{L^{4}}+A_{2}^{2}\Big{)}}-2\varepsilon,$
and thus
$\alpha^{-}(t)\leq{\frac{6}{d}\Big{(}\gamma^{2}\|v(t)\|_{L^{4}}^{4}+\gamma^{2}\|\tilde{u}(t)\|^{4}_{L^{4}}+A_{2}^{2}\Big{)}}+2\varepsilon,$
which by Lemma 2.2 provides 3.1. Now that every assumption of Lemma 3.1 holds, we can conclude that
$\lim_{t\to\infty}\|Q_{M}\xi(t)\|_{L^{2}}^{2}+\lim_{t\to\infty}\|Q_{N}\eta(t)\|_{L^{2}}^{2}=0.$
This completes the proof of Lemma 3.2. ∎
Lemma 3.3.
We set $A_{2}\coloneqq|D|^{1/2}\sum_{j=1}^{2}(|a_{j}|+|b_{j}|)$. If there exist $T>0$ and $M,N\in\mathbb{N}$ such that
(3.9) $\frac{A_{2}^{2}}{d^{2}}+\frac{\gamma^{2}}{d^{2}}\bigg{(}\limsup_{t\to\infty}\frac{1}{T}\int_{t}^{t+T}\Big{(}\|v(s)\|_{L^{4}}^{4}+\|\tilde{u}(s)\|^{4}_{L^{4}}\Big{)}\,\mathrm{d}s\bigg{)}<(\lambda_{M+1}+\lambda_{N+1}),$
then the first $(M,N)$ modes are determining for the solution $(u,v)$ to the system 1.1.
Proof.
We fix $T,t>0$ and $M,N\in\mathbb{N}$. We recall that
$K_{M,N}(t)\coloneqq{}$ ${-\gamma\big{\langle}Q_{M}\xi(t)\big{(}a_{1}-v^{2}(t)\big{)}+Q_{N}\eta(t)\big{(}b_{1}-(v(t)+\tilde{v}(t))\tilde{u}(t)\big{)},Q_{M}\xi(t)\big{\rangle}}$
$-{\gamma\big{\langle}Q_{M}\xi(t)\big{(}a_{2}+v^{2}(t)\big{)}+Q_{N}\eta(t)\big{(}b_{2}+(v(t)+\tilde{v}(t))\tilde{u}(t)\big{)},Q_{N}\eta(t)\big{\rangle}}.$
We denote $A_{2}\coloneqq|D|^{1/2}\sum_{j=1}^{2}(|a_{j}|+|b_{j}|)$ and
$H_{M,N}(\xi(s),\eta(s))\coloneqq\frac{\|\nabla Q_{M}\xi(s)\|_{L^{2}}+\|\nabla Q_{N}\eta(s)\|_{L^{2}}}{\|Q_{M}\xi(s)\|_{L^{2}}+\|Q_{N}\eta(s)\|_{L^{2}}}.$
Straightforward calculations provides
$\frac{1}{T}\int_{t}^{t+T}\bigg{(}2d\big{(}H_{M,N}(\xi(s),\eta(s))\big{)}^{2}+\frac{2|K_{M,N}(s)|}{\big{(}\|Q_{M}\xi(s)\|_{L^{2}}+\|Q_{N}\eta(s)\|_{L^{2}}\big{)}^{2}}\bigg{)}\,\mathrm{d}s$
$\geq\frac{2d}{T}\int_{t}^{t+T}\big{(}H_{M,N}(\xi(s),\eta(s))\big{)}^{2}\,\mathrm{d}s-\frac{2}{T}\int_{t}^{t+T}\Big{(}\gamma\|v(s)\|_{L^{4}}^{2}+\gamma\|\tilde{u}(s)\|^{2}_{L^{4}}+A_{2}\Big{)}\big{(}H_{M,N}(\xi(s),\eta(s))\big{)}\,\mathrm{d}s,$
and using the Cauchy–Schwarz inequality,
$\frac{1}{T}\int_{t}^{t+T}\bigg{(}2d\Big{(}H(Q_{M}\xi(s),Q_{N}\eta(s))\Big{)}^{2}+\frac{2|K_{M,N}(s)|}{\big{(}\|Q_{M}\xi(s)\|_{L^{2}}+\|Q_{N}\eta(s)\|_{L^{2}}\big{)}^{2}}\bigg{)}\,\mathrm{d}s$
$\geq\frac{2d}{T}\int_{t}^{t+T}\big{(}H_{M,N}(\xi(s),\eta(s))\big{)}^{2}\,\mathrm{d}s-\frac{2}{T}\int_{t}^{t+T}\Big{(}\gamma\|v(s)\|_{L^{4}}^{2}+\gamma\|\tilde{u}(s)\|^{2}_{L^{4}}+A_{2}\Big{)}\big{(}H_{M,N}(\xi(s),\eta(s))\big{)}\,\mathrm{d}s$
$\geq 2d\bigg{(}\frac{1}{T}\int_{t}^{t+T}\big{(}H_{M,N}(\xi(s),\eta(s))\big{)}^{2}\,\mathrm{d}s\bigg{)}$
$-{2}\bigg{(}\frac{1}{T}\int_{t}^{t+T}\Big{(}\gamma\|v(s)\|_{L^{4}}^{2}+\gamma\|\tilde{u}(s)\|^{2}_{L^{4}}+A_{2}\Big{)}^{2}\,\mathrm{d}s\bigg{)}^{\frac{1}{2}}\bigg{(}\frac{1}{T}\int_{t}^{t+T}\big{(}H_{M,N}(\xi(s),\eta(s))\big{)}^{2}\,\mathrm{d}s\bigg{)}^{\frac{1}{2}}$
$\geq\Bigg{[}2d\bigg{(}\frac{1}{T}\int_{t}^{t+T}\big{(}H_{M,N}(\xi(s),\eta(s))\big{)}^{2}\,\mathrm{d}s\bigg{)}^{\frac{1}{2}}-{6}\bigg{(}\frac{1}{T}\int_{t}^{t+T}\Big{(}\gamma^{2}\|v(s)\|_{L^{4}}^{4}+\gamma^{2}\|\tilde{u}(s)\|^{4}_{L^{4}}+A_{2}^{2}\Big{)}\,\mathrm{d}s\bigg{)}^{\frac{1}{2}}\Bigg{]}$
×${}\times\bigg{(}\frac{1}{T}\int_{t}^{t+T}\big{(}H_{M,N}(\xi(s),\eta(s))\big{)}^{2}\,\mathrm{d}s\bigg{)}^{\frac{1}{2}}\geq\delta\times\bigg{(}\frac{1}{T}\int_{t}^{t+T}\big{(}H_{M,N}(\xi(s),\eta(s))\big{)}^{2}\,\mathrm{d}s\bigg{)}^{\frac{1}{2}}$
with
$\delta\coloneqq 2d(\lambda_{M+1}+\lambda_{N+1})^{\frac{1}{2}}-6\bigg{(}\frac{1}{T}\int_{t}^{t+T}\Big{(}\gamma^{2}\|v(s)\|_{L^{4}}^{4}+\gamma^{2}\|\tilde{u}(s)\|^{4}_{L^{4}}+A_{2}^{2}\Big{)}\,\mathrm{d}s\bigg{)}^{\frac{1}{2}}.$
If $T>0$ and $M,N\in\mathbb{N}$ are such that 3.9 holds, i.e. $\delta>0$, then it implies that
$\begin{split}\liminf_{t\to\infty}&\frac{1}{T}\int_{t}^{t+T}\bigg{(}2d\bigg{(}\frac{\|\nabla Q_{M}\xi(s)\|_{L^{2}}+\|\nabla Q_{N}\eta(s)\|_{L^{2}}}{\|Q_{M}\xi(s)\|_{L^{2}}+\|Q_{N}\eta(s)\|_{L^{2}}}\bigg{)}^{2}+\frac{2|K_{M,N}(s)|}{\big{(}\|Q_{M}\xi(s)\|_{L^{2}}+\|Q_{N}\eta(s)\|_{L^{2}}\big{)}^{2}}\bigg{)}\,\mathrm{d}s>0,\end{split}$
in which case by Lemma 3.2, the first $(M,N)$ modes are determining for the system 1.1. ∎
4. Proof of Theorem 2.4
We split the proof of Theorem 2.4 into two Lemmata, then conclude. Let $(u,v)$ be any solution to the system 1.1.
Lemma 4.1.
For all $t\geq 0$ and $T>0$, it holds that
(4.1) $\frac{1}{d^{2}}\Big{(}\limsup_{t\to\infty}\frac{1}{T}\int_{t}^{t+T}\|u(s)\|^{4}_{L^{4}}\,\mathrm{d}s\Big{)}\leq\Big{(}\frac{d}{2T}+\lambda_{1}d^{2}\Big{)}\boldsymbol{\mathrm{Gr}}^{4}.$
Proof.
We fix $t\geq 0$ and $T>0$. By the property ($P_{5})$, we have
$\frac{1}{2}\frac{\,\mathrm{d}\|u(t)\|_{L^{2}}^{2}}{\,\mathrm{d}t}+d_{1}\|\nabla u(t)\|^{2}_{L^{2}}=\int_{D}(a_{1}u(t,x)+b_{1}v(t,x)-\gamma u(t,x)v^{2}(t,x)+g_{1}(t,x))u(t,x)\,\mathrm{d}x$
$\leq\int_{D}c_{1}(1-c_{2}u(t,x))u(t,x)\,\mathrm{d}x\leq c_{1}\big{(}\|u(t)\|_{L^{2}}-c_{2}\|u(t)\|_{L^{2}}^{2}\big{)}\leq B_{2}\|u(t)\|_{L^{2}},$
Where $B_{2}\coloneqq c_{1}|D|^{1/2}$.from which, on one hand, we obtain
$\frac{1}{2}\frac{\,\mathrm{d}\|u(t)\|_{L^{2}}^{2}}{\,\mathrm{d}t}+d_{1}\|\nabla u(t)\|^{2}_{L^{2}}\leq B_{2}\|u(t)\|_{L^{2}},$
and, therefore,
$\limsup_{t\to\infty}\|u(t)\|_{L^{2}}\leq\frac{B_{2}}{\lambda_{1}d_{1}}\leq d\Big{(}\frac{F}{\lambda_{1}d^{2}}\Big{)}=d\boldsymbol{\mathrm{Gr}}.$
On the other hand, we have also for every $T\geq 0$
$\frac{1}{2}\|u(t+T)\|^{2}_{L^{2}}+d_{1}\int_{t}^{t+T}\|\nabla u(s)\|^{2}_{L^{2}}\,\mathrm{d}s\leq\frac{1}{2}\|u(t)\|_{L^{2}}^{2}+B_{2}\int_{t}^{t+T}\|u(s)\|_{L^{2}}\,\mathrm{d}s.$
It follows that
$\limsup_{t\to\infty}\frac{1}{T}\int_{t}^{t+T}\|\nabla u(s)\|^{2}_{L^{2}}\,\mathrm{d}s\leq\frac{d\boldsymbol{\mathrm{Gr}}^{2}}{2T}+B_{2}\boldsymbol{\mathrm{Gr}}=\frac{d\boldsymbol{\mathrm{Gr}}^{2}}{2T}+\lambda_{1}d^{2}\big{(}\frac{B_{2}}{\lambda_{1}d^{2}}\big{)}\boldsymbol{\mathrm{Gr}}=\Big{(}\frac{d}{2T}+\lambda_{1}d^{2}\Big{)}\boldsymbol{\mathrm{Gr}}^{2}.$
We recall, see [2, Proposition III.2.35], that
$\|u(s)\|^{4}_{L^{4}}\leq\|u(s)\|^{2}_{L^{2}}\|\nabla u(s)\|^{2}_{L^{2}}.$
Thus, we have
$\Big{(}\limsup_{t\to\infty}\frac{1}{T}\int_{t}^{t+T}\|u(s)\|^{4}_{L^{4}}\,\mathrm{d}s\Big{)}$ $\leq\Big{(}\limsup_{t\to\infty}\frac{1}{T}\int_{t}^{t+T}\|u(s)\|^{2}_{L^{2}}\|\nabla u(s)\|^{2}_{L^{2}}\,\mathrm{d}s\Big{)}$
$\leq\Big{(}\limsup_{t\to\infty}\|u(t)\|_{L^{2}}\Big{)}^{2}\Big{(}\limsup_{t\to\infty}\frac{1}{T}\int_{t}^{t+T}\|\nabla u(s)\|^{2}_{L^{2}}\,\mathrm{d}s\Big{)}.$
It implies that
$\frac{1}{d^{2}}\Big{(}\limsup_{t\to\infty}\frac{1}{T}\int_{t}^{t+T}\|u(s)\|^{4}_{L^{4}}\,\mathrm{d}s\Big{)}\leq\Big{(}\frac{d}{2T}+\lambda_{1}d^{2}\Big{)}\boldsymbol{\mathrm{Gr}}^{4},$
and, hence we get 4.1. ∎
Lemma 4.2.
For all $t\geq 0$ and $T>0$, it holds that
(4.2) $\frac{1}{d^{2}}\Big{(}\limsup_{t\to\infty}\frac{1}{T}\int_{t}^{t+T}\|v(s)\|^{4}_{L^{4}}\,\mathrm{d}s\Big{)}\leq\Big{(}\frac{d}{2T}+\lambda_{1}d^{2}\Big{)}\boldsymbol{\mathrm{Gr}}^{4}.$
Proof.
We fix $t\geq 0$, $T>0$, and $x\in D$. Observe that $w\coloneqq u+v$ satisfies to
$\partial_{t}w(t,x)=d_{1}\Delta u(t,x)+d_{2}\Delta v(t,x)+(a_{1}+a_{2})u(t,x)+(b_{1}+b_{2})v(t,x)+(g_{1}+g_{2})(t,x).$
Here, we use the property ($P_{6})$ and recall that $u,v\geq 0$. Then, we have
$\frac{1}{2}\frac{\,\mathrm{d}\|w(t)\|_{L^{2}}^{2}}{\,\mathrm{d}t}+d_{1}\langle\nabla u(t),\nabla w(t)\rangle+d_{2}\langle\nabla v(t),\nabla w(t)\rangle=\langle(g_{1}+g_{2})(t),w(t)\rangle\leq\|(g_{1}+g_{2})(t)\|_{L^{2}}\|w(t)\|_{L^{2}}$
from which we obtain
$\frac{\,\mathrm{d}\|w(t)\|_{L^{2}}}{\,\mathrm{d}t}+d\lambda_{1}\|w(t)\|_{L^{2}}\leq\|(g_{1}+g_{2})(t)\|_{L^{2}}$
and
(4.3) $\limsup_{t\to\infty}\|w(t)\|_{L^{2}}\leq d\Big{(}\frac{g_{1}^{*}+g_{2}^{*}}{\lambda_{1}d^{2}}\Big{)}\leq d\Big{(}\frac{F}{\lambda_{1}d^{2}}\Big{)}=d\boldsymbol{\mathrm{Gr}}.$
We have also
$\frac{1}{2}\|w(t+T)\|^{2}_{L^{2}}+d\int_{t}^{t+T}\|\nabla w(s)\|_{L^{2}}^{2}\,\mathrm{d}s\leq\frac{1}{2}\|w(s)\|_{L^{2}}^{2}$
$+\int_{t}^{t+T}\|g_{1}(s)\|_{L^{2}}\|w(s)\|_{L^{2}}\,\mathrm{d}s+\int_{t}^{t+T}\|g_{2}(s)\|_{L^{2}}\|w(s)\|_{L^{2}}\,\mathrm{d}s,$
and using 4.3, it follows
$\limsup_{t\to\infty}\frac{1}{T}\int_{t}^{t+T}\|\nabla w(s)\|_{L^{2}}^{2}\,\mathrm{d}s\leq\frac{d}{2T}\boldsymbol{\mathrm{Gr}}^{2}+(g_{1}^{*}+g_{2}^{*})\boldsymbol{\mathrm{Gr}}$
$\leq\frac{d}{2T}\boldsymbol{\mathrm{Gr}}^{2}+\lambda_{1}d^{2}\frac{(g^{*}+h^{*})}{\lambda_{1}d^{2}}\boldsymbol{\mathrm{Gr}}\leq\Big{(}\frac{d}{2T}+\lambda_{1}d^{2}\Big{)}\boldsymbol{\mathrm{Gr}}^{2}.$
Upon recalling that
$\Big{(}\limsup_{t\to\infty}\frac{1}{T}\int_{t}^{t+T}\|w(s)\|^{4}_{L^{4}}\,\mathrm{d}s\Big{)}\leq\Big{(}\limsup_{t\to\infty}\|w(t)\|_{L^{2}}\Big{)}^{2}\Big{(}\limsup_{t\to\infty}\frac{1}{T}\int_{t}^{t+T}\|\nabla w(s)\|_{L^{2}}^{2}\,\mathrm{d}s\Big{)},$
and that $w=u+v$, we obtain
$\frac{1}{d^{2}}\Big{(}\limsup_{t\to\infty}\frac{1}{T}\int_{t}^{t+T}\|v(s)\|^{4}_{L^{4}}\,\mathrm{d}s\Big{)}\leq\Big{(}\frac{d}{2T}+\lambda_{1}d^{2}\Big{)}\boldsymbol{\mathrm{Gr}}^{4},$
which completes the proof of 4.2. ∎
Proof of Theorem 2.4
We fix $t\geq 0$ and $T>0$. To complete the proof of Theorem 2.4, we combine 4.14.2 to get
$\frac{A_{2}^{2}}{d^{2}}+\frac{\gamma^{2}}{d^{2}}\Big{(}\limsup_{t\to\infty}\frac{1}{T}\int_{t}^{t+T}\big{(}\|u(s)\|^{4}_{L^{4}}+\|v(s)\|^{4}_{L^{4}}\big{)}\,\mathrm{d}s\Big{)}$
$\leq\gamma^{2}\Big{(}\frac{d}{T}+2\lambda_{1}d^{2}\Big{)}\boldsymbol{\mathrm{Gr}}^{4}+\frac{A_{2}^{2}}{d^{2}}\leq\gamma^{2}\Big{(}\frac{d}{T}+2\lambda_{1}d^{2}\Big{)}\boldsymbol{\mathrm{Gr}}^{4}+\lambda_{1}\boldsymbol{\mathrm{Gr}}.$
Let $M,N\in\mathbb{N}$ such that
$(\lambda_{M+1}+\lambda_{N+1})>2\gamma^{2}\lambda_{1}d^{2}\boldsymbol{\mathrm{Gr}}^{4}+\lambda_{1}\boldsymbol{\mathrm{Gr}},$
then it follows that
$\frac{A_{2}^{2}}{d^{2}}+\frac{\gamma^{2}}{d^{2}}\Big{(}\limsup_{t\to\infty}\frac{1}{T}\int_{t}^{t+T}\|u(s)\|^{4}_{L^{4}}+\|v(s)\|^{4}_{L^{4}}\,\mathrm{d}s\Big{)}<(\lambda_{M+1}+\lambda_{N+1}),$
for $T$ large enough, i.e.
$T>\frac{d\boldsymbol{\mathrm{Gr}}^{4}}{(\lambda_{M+1}+\lambda_{N+1})-2\gamma^{2}\lambda_{1}d^{2}\boldsymbol{\mathrm{Gr}}^{4}-\lambda_{1}\boldsymbol{\mathrm{Gr}}}.$
Thus, by Lemma 3.3, the first $(M,N)$ modes are determining for the couple of solution $(u,v)$ to the system 1.1. This completes the proof of the main result Theorem 2.4. ∎
Acknowledgement
This work was supported by the Austrian Science Fund (FWF) 10.55776/P44681.